The following data are ages at death in years for a sample of people who were all born in the same year: 11, 34, 42, 44, 48, 50, 54, 55,55,58, 59, 60, 65, 67, 68, 68, 69, 70, 70, 72, 74, 76, 76, 79, 81, 83, 85, 86. What is the median age at death for these data?
Answer #1

Solution :


11,34,42,44,48,50,54,55,55,58,59,60,65,67,68,68,69,70,70,72,74,76,76,79,81,83,85,86

Here, n=28

Q2 = (2(n+1)/4)th value of the observation

= (2 * 29 / 4)th value of the observation

= (14.5)th value of the observation

= 14th observation + 0.5[15th - 14th]

= 67 + 0.5[68 - 67]

= 67 + 0.5(1)

= 67 + 0.5

= 67.5

.Median = 67.5